# Questions: A Taxonomy for Critical Reflection in Machine-Supported Decision-Making

### v1

@@ Abstract.
Decision-makers run the risk of relying too much on machine recommendations. Explainable AI, a common strategy for calibrating reliance, has mixed and even negative effects, such as increasing overreliance. To cognitively engage the decision-maker and to facilitate a deliberate decision-making process, we propose a potential ‘reflection machine’ that supports critical reflection about the pending decision, including the machine recommendation. Reflection has been shown to improve critical thinking and reasoning, and thus decision-making. One way to stimulate reflection is to ask relevant questions. To systematically create questions, we present a question taxonomy inspired by Socratic questions and human-centred explainable AI. This taxonomy can contribute to the design of such a ‘reflection machine’ that asks decision-makers questions. Our work is part of the growing research on human-machine collaborations that goes beyond the paradigm of machine recommendations and explanations, and aims to enable greater human oversight as required by the European AI Act.
Reflective Practice, Socratic Questions, Cognitive Intervention, Human-AI Collaboration, Decision-Support Systems, Appropriate Reliance
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: ACM Conference; 978-1-4503-XXXX-X/18/06isbn: ;ccs: Human-centered computing Collaborative interactionccs: Computing methodologies Artificial intelligenceccs: Human-centered computing Interaction design theory, concepts and paradigmsccs: Human-centered computing Computer supported cooperative work
@@ 1. Introduction
Reflective thinking is an important skill in professional practices (Schön, 1983). It is beneficial in, for example, management (Cunliffe, 2002), law (Casey, 2014), and healthcare (Mann et al., 2009; Sandars, 2009), and thus part of many educational curricula (Boud and Walker, 1998). Reflection, as John Loughran (1996, p14) states, is a “deliberate and purposeful act”, which, according to John Dewey (1933, p9), consists of “active, persistent and careful consideration”. During decision-making, reflection has the potential to improve reasoning, judgement, and problem-solving (Ghanizadeh, 2017; Khoshgoftar and Barkhordari-Sharifabad, 2023; Mamede and Schmidt, 2023), as it allows to evaluate the validity of information and assumptions (Mezirow, 1990). Consequently, reflection has been shown to improve strategic decisions (Walger et al., 2016) and diagnostic accuracy (Mamede et al., 2008; Prakash et al., 2019; Hess et al., 2015). In addition, reflection enables decision-makers to be more aware of the reasons for their decisions, which supports them in taking responsibility and being accountable. Ideally, responsible decisions lead to more desirable and fairer outcomes for all parties involved. In view of these effects, it seems sensible to encourage and support reflection during decision-making. Especially given the current paradigm in which many high-stakes decisions are supported by machine recommendations.
@@
Decision support systems (DSS) assist decision-makers in domains like law, finance and healthcare. Physicians, for example, can use clinical decision support systems to find a treatment option for a patient. Studies show, however, that operators tend to rely too much on these systems and accept incorrect recommendations (Dratsch et al., 2023), which is known as overreliance (Passi and Vorvoreanu, 2022). A common approach to mitigating overreliance is explainable AI (XAI) (Longo et al., 2024), which focuses on providing information about which input was decisive for a particular output. Provided explanations are however often not taken into account (van der Waa et al., 2021), or they can reinforce the acceptance of any machine recommendation regardless of its quality or correctness (Bansal et al., 2021; Bussone et al., 2015; Ehsan et al., 2021). In light of this, and to reduce harmful consequences of overreliance, policymakers and laws such as the European AI Act require more significant, i.e., effective, human oversight over algorithmic decisions, enabling decision-makers to adequately scrutinise machine recommendations before making a decision (Pariliament, 2024).
In this paper, we propose a potential solution for mitigating overreliance, namely supporting the decision-maker in critically reflecting on the pending decision, including the machine recommendation. By critical reflection, we mean the practice of evaluating relevant information, without taking information at face value, and scrutinising (taken for granted) assumptions (Hong and Choi, 2011; Sengers et al., 2005). In relation to a DSS, this means that the decision-maker should remain critical of the output and data, as well as be open to alternative approaches. As mentioned, reflection can increase reasoning and allows to identify the reasons for a decision (Ghanizadeh, 2017). We therefore hypothesise that critical reflection can increase cognitive engagement of the decision-maker and thus help to reduce over-reliance on decision-support systems.
@@
In order to elicit reflection, the decision-maker can be asked questions, particularly in the form of the Socratic questioning method. The Socratic method, named after the Greek philosopher Socrates, involves systematically asking questions in order to explore complex ideas, to uncover and scrutinise assumptions, and to gain a better understanding of the topic under investigation (Paul, 2019). Although the Socratic method is slowly gaining traction in the field of human-machine collaboration (Liu et al., 2024a, b; Lara, 2021; Ang et al., 2023), there is no systematic approach as to how this questioning method can be applied to the the context of machine-supported decision-making. We therefore provide a taxonomy of questions for critical reflection by synthesising and combining prior work, namely 1) a taxonomy of Socratic questions (Paul, 2019), 2) a question bank for the design of human-centred XAI (Liao et al., 2020, 2021; Liao and Varshney, 2022), and 3) a revised version of Bloom’s taxonomy for categorising educational objectives (Krathwohl, 2002; Forehand, 2010). In doing so, this paper provides an answer to the question: How to create relevant questions that stimulate critical reflection in the context machine-supported decision-making? Although our question taxonomy is applicable to any decision-making domain, we primarily provide examples from the area of clinical decision-making, where decisions are often complex and have major implications.
This paper makes the following contributions:
  • We make the case for a potential ‘reflection machine’ that supports critical reflection through questions, by, among other things, repurposing explanations (section 2.2), which functions as a potential solution to reduce overreliance on DSS (section 2.3).
@@
  • We discuss literature that shows that reflection increases cognitive engagement and reasoning, and thus decision-making, and that questions can help stimulate reflection (section 2.4), in particular in the form of Socratic questions (section 2.5).
  • We propose a new question taxonomy to help create relevant questions to stimulate reflection during machine-supported decision-making (section 4). In doing so, we concretise current conceptual proposals for human-machine collaborations that go beyond machine recommendations and explanations. This taxonomy provides designers and developers with the foundations for designing a ‘reflection machine’ that asks questions to the decision-maker.
@@ 2. Background
In the following section we will briefly discuss background work that led us to create our question taxonomy. In order to address the problem of overreliance on machine recommendations and explanations (section 2.1), we begin with a short review of the relevant literature on human-machine collaboration, particularly in relation to the human factor in explainable AI (section 2.2), and alternative strategies to reduce overreliance (section 2.3). From this, we derive that cognitive engagement is a promising approach. We identify the benefits of reflective thought for cognitive engagement by promoting critical thinking, reasoning, and thus responsible decision-making (section 2.4). Finally, we suggest that questions can be useful to stimulate reflection. We focus on Socratic questions, which are widely used in education (section 2.5). In short, to arrive at our taxonomy, we combine relevant insights from empirical findings and literature on human-machine collaboration, education, and philosophy.
2.1. Overreliance in Human-Machine Decision-Making
@@
A meta-study of 106 experimental studies on human-machine collaboration found that decision-making performance generally decreases when decision-makers use decision-support systems (DSS) (Vaccaro et al., 2024). Decision-makers may tend to prefer machine recommendations over their own judgement (Tschandl et al., 2020), even when the recommendations are wrong (Jacobs et al., 2021; van der Stigchel et al., 2023). The acceptance of incorrect recommendations is referred to as overreliance (Passi and Vorvoreanu, 2022). As briefly mentioned in the introduction, a prominent method to calibrate reliance on DSS is explainable AI (XAI) (Vasconcelos et al., 2023). It is assumed that insight into how a DSS computed a decision will help the operator to assess the machine recommendation and thus make an informed decision about whether or not to follow the recommendation. The aforementioned meta-study on human-machine collaboration found that explanations do not lead to significant synergy effects in terms of the performance of the human-machine team (Vaccaro et al., 2024). On the contrary, explanations may increase the likelihood of operators accepting a machine recommendation regardless of its correctness (Bansal et al., 2021; Bussone et al., 2015; Ehsan et al., 2021), thereby increasing overreliance (Vered et al., 2023). One reason for this is that the focus on numerical explanations can give a false sense of objectivity, leading operators, regardless of their expertise, to overestimate the capabilities of a DSS (Ehsan et al., 2024). As such, the provision of additional information in the form of explanations does not automatically lead to more cognitive engagement of the decision-maker (Gajos and Mamykina, 2022).
2.2. Human-Centred XAI
Work is being done on more effective explanations by moving from the current algorithmic focus of XAI methods to a broader, socio-technical perspective of explanations (Liao and Varshney, 2022; Ehsan et al., 2021; Ehsan and Riedl, 2024; Ghassemi et al., 2021). Given the diverse needs of different stakeholders in the human-machine collaboration ecosystem, e.g., as debugging models, assessing regulatory compliance, making informed decisions, or contesting automated decision, current explainable AI methods that focus on providing a technical answer to how a DSS computed a particular output may not be always appropriate (Chari et al., 2024; Liao and Varshney, 2022; Brennen, 2020). As one participant working on clinical decision-support systems mentioned in a study by Liao et al. (2020) on human-centred explanations:
@@ 2.2. Human-Centred XAI
“[explanations by system rationale] are essentially ‘this is how I do it, take it or leave it’. But doctors don’t like this approach…Thinking that [AI is] giving treatment recommendations is the wrong place to start, because doctors know how to do it. It’s everything that happens around that decision they need help with… more discussions about the output, rather than how you get there” (p6).
So even if it is known how a DSS computed an output, operators might be more interested in contextual factors (Zednik, 2019; Ehsan and Riedl, 2024), which relate to questions like, “Is the data representative of the current situation?”, “Is the data up-to-date?”, “Are there any other overlooked possibilities?”. To stimulate “more discussions about the output” (study participant in Liao et al., 2020), we want to utilise current explainable AI methods and repurpose explanations to help formulate and inform questions that can be asked of decision-makers. As such, we consider “explanations as a means to help a data subject [and decision-maker] act rather than merely understand” (Wachter et al., 2017, p843), where ‘act’ in this case is deliberate decision-making.
2.3. Alternative Approaches to Reduce Overreliance
@@
To address the remaining challenge of overreliance, scholars explore different interactions with machine recommendations and explanations that attempt to cognitively engage the decision-maker in the decision-making process. We will briefly mention three relevant approaches: 1) promoting cognitive engagement through interventions, 2) supporting the decision-maker to make their own decision by presenting evidence for and against, and 3) formulating explanations as questions.
@@
First, cognitive interventions are based on the dual-process model of reasoning, and aim to interrupt the decision-maker’s habitual thinking (system 1) in order to encourage analytical thinking (system 2) (Lambe et al., 2016). Cognitive interventions can take the form of checklists, instructions for analytical thinking, or reflection, for example (Lambe et al., 2016; Prakash et al., 2019). In the context of human-machine collaboration, one study tested a cognitive intervention strategy by presenting explanations at different times, such as delayed or on-demand. These interventions were found to reduce, yet not eliminate, overreliance on DSS compared to the direct presentation of explanations (Buçinca et al., 2021).
@@
Second, instead of presenting a recommendation and justifying it with an explanation that the decision-maker must accept or reject, a so-called hypothesis-driven recommender shows evidence for and against a decision (Miller, 2023). In this way, the system supports the decision-maker’s cognitive process and gives them the control to make an informed decision themselves. Another study found that while presenting evidence for and against did little to improve diagnostic accuracy, physicians valued the reflective aspect of it and had more confidence in their final decisions (Cabitza et al., 2023).
@@
Third, in an experiment, causal explanations were presented both as statements and questions, and compared with each other (Danry et al., 2023). An explanation as a statement is, for example, “The Diagnosis $Y$ is because of symptom $x$”, and formulated as question: “If the patient has symptom $x$, does it necessarily mean they have condition $Y$?”. The study found that explanations framed as questions improved human judgement about the logical validity of the information provided. Similar to the hypothesis-driven recommender, questions help the decision-maker to think for themselves (Ryan and Deci, 2000), leading to more cognitive engagement (Gajos and Mamykina, 2022).
@@
In addition to these approaches, it has been suggested that DSS should take on roles other than that of a recommender, such as an analyser presenting counterarguments, or as a devil’s advocate raising objections and challenging the decision-maker (Chiang et al., 2024; Sarkar, 2024; Ma et al., 2024a; Miller, 2023). Insights from decision-making in human groups show that these different roles, i.e., devil’s advocate and analyser, have positive effects on decision performance (Schwenk, 1984; Schweiger et al., 1989). Further, other authors argue that DSS should support deliberation and facilitate discussion between human and machine (Ma et al., 2024b; Zhang et al., 2023; Haselager et al., 2023; Schmidt and Mamede, 2023). Similarly, we imagine a ‘reflection machine’ that functions as a cognitive intervention by posing questions to the decision-maker to encourage them to reflect on the pending decision, including machine recommendations. Our question taxonomy for creating relevant, thought-provoking questions forms the foundation for the design of such a ‘reflection machine’.

### v2

# A Taxonomy of Questions for Critical Reflection in Machine-Assisted Decision-Making

@@ Abstract.
Decision-makers run the risk of relying too much on machine recommendations, which is associated with lower cognitive engagement. Reflection has been shown to increase cognitive engagement and improve critical thinking and reasoning and therefore decision-making. However, there is currently no approach to support reflection in machine-assisted decision-making. We therefore present a taxonomy that serves to systematically create questions related to machine-assisted decision-making that promote reflection and thus cognitive engagement and ultimately a deliberate decision-making process. Our taxonomy builds on a taxonomy of Socratic questions and a question bank for human-centred explainable AI (XAI), and illustrates how XAI techniques can be utilised and repurposed to formulate questions. As a use case, we focus on clinical decision-making. An evaluation in education confirms the applicability and expected benefits of our taxonomy. Our work contributes to the growing research on human-AI interaction that goes beyond the paradigm of machine recommendations and explanations and aims to enable effective human oversight as required by the European AI Act.
Reflection, Cognitive Engagement, Decision-Support Systems, Overreliance, Critical Thinking, Socratic Questioning, Human-AI Interaction
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: ; isbn: ;ccs: Human-centered computing Collaborative interactionccs: Computing methodologies Artificial intelligenceccs: Human-centered computing Interaction design theory, concepts and paradigmsccs: Human-centered computing Computer supported cooperative work
@@ 1. Introduction
Reflection, also referred to as reflective practice or reflective thought, is an important skill in professional practices (Schön, 1983). It is beneficial in, for example, management (Cunliffe, 2002), law (Casey, 2014), and healthcare (Mann et al., 2009; Sandars, 2009), and is part of many educational curricula (Boud and Walker, 1998). As the American philosopher John Dewey (1933, p.9) states, reflection consists of an
“active, persistent and careful consideration of any belief or supposed form of knowledge in light of the grounds that support it and further conclusion to which it tends”.
@@
As reflection helps to scrutinise the validity of information and assumptions (Mezirow, 1990), it has the potential to improve reasoning, judgement, and problem-solving (Ghanizadeh, 2017; Khoshgoftar and Barkhordari-Sharifabad, 2023; Mamede and Schmidt, 2023). Consequently, reflection has been shown to improve strategic decisions (Walger et al., 2016) and diagnostic accuracy (Mamede et al., 2008; Prakash et al., 2019; Hess et al., 2015). In addition, reflection enables decision-makers to recognise and articulate the reasons for their decisions, which helps them to direct their actions accordingly and take responsibility (forward-looking) and accountability (backward-looking). Responsible decision-making ideally leads to more desirable and fairer outcomes for all parties involved. In view of these effects, it seems sensible to encourage and support reflection during decision-making, especially as many high-impact decisions nowadays are informed by machine recommendations.
@@
Decision-support systems (DSS) assist decision-makers in domains like law, finance and healthcare. Physicians, for example, can use clinical DSSs to make a diagnosis or find a treatment option for a patient. Studies show, however, that operators tend to rely too much on these systems and accept incorrect recommendations (Dratsch et al., 2023), which is known as overreliance (Passi and Vorvoreanu, 2022). Moreover, explainable AI (XAI) (Longo et al., 2024), a common approach for calibrating reliance, has mixed and even negative effects, such as increasing overreliance (Bansal et al., 2021; Bussone et al., 2015; Ehsan et al., 2021). In light of this, and to reduce harmful consequences of overreliance, the European AI Act (European Parliament and Council of the European Union, 2024, §14) requires:
[…] that natural persons to whom human oversight is assigned are enabled, as appropriate and proportionate: […] (b) to remain aware of the possible tendency of automatically relying or over-relying on the output produced by a high-risk AI system (automation bias), in particular for high-risk AI systems used to provide information or recommendations for decisions to be taken by natural persons;
In this paper, we propose a potential solution for mitigating overreliance, namely cognitive engagement of the decision-maker through questions that encourage critical reflection on the decision at hand, including the machine recommendation and the data that led to it. By critical reflection, we mean the above-mentioned practice of evaluating relevant information, without taking information at face value, and analysing assumptions that may have been taken for granted (Hong and Choi, 2011; Sengers et al., 2005). In order to elicit reflection, the decision-maker can be asked questions, for which the Socratic questioning method offers a helpful starting point. The Socratic method, named after the Greek philosopher Socrates, involves systematically asking questions to explore complex ideas, uncover assumptions, and gain a better understanding of the topic under investigation (Paul, 2019).
@@
Although the Socratic method is slowly gaining traction in the field of human-machine interaction (Liu et al., 2024; Lara, 2021; Ang et al., 2023), there is as yet no systematic approach to applying this questioning method to the context of machine-assisted decision-making. To address this gap, we therefore provide a taxonomy of questions that serves the creation of questions for critical reflection in machine-assisted decision-making. For our taxonomy we synthesise prior work, namely 1) a taxonomy of Socratic questions (primarily for education) (Paul, 2019), 2) a question bank for the design of human-centred XAI (Liao et al., 2020, 2021; Liao and Varshney, 2022), and 3) a revised version of Bloom’s taxonomy for categorising educational objectives (Krathwohl, 2002; Forehand, 2010).
We primarily focus on clinical decision-making, where a physician uses a DSS to find a treatment option for patients. In this context, there may be multiple solutions with uncertain outcomes, so deliberate decision-making is crucial. An evaluation of the applicability of our taxonomy in the education domain shows however that our taxonomy can be applied to other domains without adaptation. Overall, our proposed taxonomy builds on evidence from literature that critical reflection can increase cognitive engagement of the decision-maker. This paper makes the following contributions:
  • With respect to overreliance (section 2.1) and human-centred XAI (section 2.2), we add to the growing body of literature on approaches to increase cognitive engagement (section 2.3) and, in particular, work towards and concretise existing proposals to stimulate reflection.
@@
  • We discuss literature that shows that reflection increases cognitive engagement and reasoning, and thus decision-making (section 2.4), and that questions can help stimulate reflection (section 2.5).
  • We present a new taxonomy to help formulate relevant questions that stimulate reflection in machine-assisted decision-making (section 4). This taxonomy supports the design and development of interventions that go beyond machine recommendations and explanations. In addition, our taxonomy can benefit other stakeholders such as developers, auditors, or consumers who need to remain critical in the development, evaluation, and use of DSS (section 6).
@@ 2. Background
In the following, we will discuss background work that led us to the creation of our question taxonomy. In order to sketch the problem of overreliance on machine recommendations and explanations (section 2.1), we begin with a short review of the relevant literature on human-machine interaction, particularly in relation to the human factor in explainable AI (section 2.2). From alternative interaction approaches, we derive that encouraging cognitive engagement is a promising method to reduce overreliance (section 2.3). We identify the benefits of reflective thought for cognitive engagement as it promotes critical thinking, reasoning, and thus responsible decision-making (section 2.4). Finally, we suggest that questions can be useful to stimulate reflection and thus cognitive engagement, and introduce the Socratic questioning method, which is widely used in education (section 2.5). In short, to arrive at our taxonomy, we combine relevant insights from empirical findings and literature on human-machine interaction, education, and philosophy.
2.1. Overreliance on DSS
Decision-support systems (DSS) generally process information to arrive at an output, i.e., recommendation, by using (simple) mathematical models, or more advanced machine learning models, including deep learning approaches. A meta-study of 106 experimental studies on machine-assisted decision-making found that decision-making accuracy generally decreases when decision-makers use DSS (Vaccaro et al., 2024). Decision-makers may tend to prefer machine recommendations over their own judgement (Tschandl et al., 2020), even when the recommendations are wrong (Jacobs et al., 2021; van der Stigchel et al., 2023). The acceptance of incorrect recommendations is referred to as overreliance, which can be because of low vigilance or lack of cognitive engagement (Passi and Vorvoreanu, 2022).
@@ 2.1. Overreliance on DSS
A common approach to mitigating overreliance is explainable AI (XAI) (Vasconcelos et al., 2023), which can, for example, provide information about which input was decisive for a particular output, i.e., recommendation. It is assumed that these explanations will help the operator to assess the machine recommendation and thus make an informed decision about whether or not to follow the recommendation. Provided explanations are, however, not always taken into account (van der Waa et al., 2021). Further, the aforementioned meta-study also found that explanations do not lead to significant synergy effects in terms of the performance of the human-machine team (Vaccaro et al., 2024). On the contrary, explanations may increase the likelihood of operators accepting a machine recommendation regardless of its correctness (Bansal et al., 2021; Bussone et al., 2015; Ehsan et al., 2021), thereby increasing overreliance (Vered et al., 2023). One reason for this is that the focus on numerical explanations can give a false sense of objectivity, leading operators, regardless of their expertise, to overestimate the capabilities of a DSS (Ehsan et al., 2024). Consequently, the provision of additional information in the form of explanations does not automatically lead to more cognitive engagement of the decision-maker and more accurate decisions (Gajos and Mamykina, 2022).
2.2. A Question Bank for XAI
To provide more actionable explanations based on user needs, a question bank for explainable AI helps designers and developers select the appropriate XAI technique to provide a relevant explanation that supports the operator’s goal, e.g., informed decision-making. To this end, the question bank maps prototypical user questions to available XAI techniques (Liao et al., 2020, 2021; Liao and Varshney, 2022). Feature importance, for example, can provide insights into why-questions, such as “Why was this prediction given?”.
Answering the (technical) question of how a DSS arrived at a particular output is however not always desired or helpful. As one participant working on clinical decision-support systems mentioned in the study on the XAI question bank:
@@ 2.2. A Question Bank for XAI
“[explanations by system rationale] are essentially ‘this is how I do it, take it or leave it’. But doctors don’t like this approach…Thinking that [AI is] giving treatment recommendations is the wrong place to start, because doctors know how to do it. It’s everything that happens around that decision they need help with… more discussions about the output, rather than how you get there” (Liao et al., 2020, p.6).
While the XAI question bank allows for the design of more human-centred XAI, it is still very algorithmic- and data-centred, as we will show (Figure 1) and as the authors themselves note. Even if it is known how a DSS arrived at an output or recommendation, operators might be more interested in factors outside the model (Zednik, 2019; Ehsan and Riedl, 2024), which could relate to questions like, “What is missing from the dataset?”, “Are there any other overlooked alternative solutions?”, or “How to integrate this recommendation into the decision-process?”. Scholars therefore advocate moving from the current algorithmic focus of XAI methods to a broader, socio-technical perspective of explanations (Ehsan et al., 2021; Ehsan and Riedl, 2024; Ghassemi et al., 2021).
With our proposed taxonomy, we want to stimulate “more discussions about the output” (study participant in Liao et al., 2020) by identifying and categorising relevant question types. As we will show, current XAI techniques can be utilised and explanations can be repurposed to formulate and inform possible questions. Framing causal explanations as questions, for example, was found to improve human judgement (Danry et al., 2023). As such, in line with actionability, we consider “explanations as a means to help a data subject [and decision-maker] act rather than merely understand” (Wachter et al., 2017, p.843), where ‘act’ in this case is deliberate decision-making through reflection.
2.3. Increasing Cognitive Engagement
@@
To mitigate overreliance, scholars explore different interactions with machine recommendations and explanations that attempt to cognitively engage the decision-maker in the decision-making process. We will briefly mention four relevant approaches: 1) promoting cognitive engagement through interventions, 2) supporting the decision-maker to make their own decision by presenting evidence for and against, 3) formulating explanations as questions, and 4) proposals to stimulate reflection.
@@
First, cognitive interventions are based on the dual-process model of reasoning, and aim to interrupt the decision-maker’s habitual thinking (system 1) in order to encourage analytical thinking (system 2) (Lambe et al., 2016). Cognitive interventions can take the form of checklists, instructions for analytical thinking, or reflection, for example (Lambe et al., 2016; Prakash et al., 2019). In the context of human-machine interaction, one study found that such interventions, i.e., presenting explanations delayed, on-demand, or only after the operator made a decision, can reduce, yet not eliminate, overreliance on DSS (Buçinca et al., 2021).
@@
Second, instead of presenting a recommendation and justifying it with an explanation that the decision-maker must accept or reject, a so-called hypothesis-driven recommender shows evidence for and against a decision (Miller, 2023). In this way, the system supports the decision-maker’s reasoning process and gives them the control to make an informed decision themselves. A study found that while presenting evidence for and against did little to improve diagnostic accuracy, physicians valued the reflective aspect and had more confidence in their final decisions (Cabitza et al., 2023).
@@
Third, in an experiment causal explanations were presented both as statements and questions (Danry et al., 2023). For example, the explanation as a statement: “The patient has the flu because of a headache”, is formulated as a question: “If the patient has a headache, does it follow that they have the flu?”. The study found that explanations framed as questions improved human judgement about the logical validity of the information provided. Similar to the hypothesis-driven recommender, questions help the decision-maker to think for themselves (Ryan and Deci, 2000), which leads to more cognitive engagement (Gajos and Mamykina, 2022).
@@
Fourth, it has been suggested that DSS should take on roles other than that of a recommender, such as a devil’s advocate raising objections and challenging the decision-maker (Chiang et al., 2024; Sarkar, 2024; Ma et al., 2024b). Insights from decision-making in human groups show that a devil’s advocate has positive effects on decision performance (Schwenk, 1984; Schweiger et al., 1989). Similarly, others argue that DSS should support deliberation and facilitate reflection (Ma et al., 2024a; Zhang et al., 2023; Haselager et al., 2023; Schmidt and Mamede, 2023; Cornelissen et al., 2022).
In view of the above, we understand questions as cognitive interventions (1), which facilitate own reasoning and decision-making (2&3). With our taxonomy, we therefore aim to both build on and work towards these four approaches by offering more potential questions compared to 3) and concretising the proposals for promoting deliberation and reflection (4).